OpenStudy (anonymous):

Is the following always, sometimes or never true? x + 1 = x + 0 A. Always B. Sometimes C. Never

OpenStudy (anonymous):

A

myininaya (myininaya):

Please guys explain your answers. x+1=x+0 is never true since 1 never equals 0

OpenStudy (anonymous):

never

OpenStudy (anonymous):

Always trust Myininaya :D

hero (hero):

x + 1 = x + 0 -x -x ____________ 1 = 0 false

OpenStudy (anonymous):

(0) + 1 = (1) + 0 answer = 1

myininaya (myininaya):

x cannot take on two values at the same time

OpenStudy (anonymous):

I see what you mean.. my mistake.